OpenStudy (anonymous):

help me find the frequency of d = 2 sin ((pi/3)t)!!

jimthompson5910 (jim_thompson5910):

What's the period of that trig equation?

OpenStudy (anonymous):

ummm... i struggle with periods as well .-.

jimthompson5910 (jim_thompson5910):

y = A*sin(Bx - C) + D is the general equation with period T = 2pi/B

jimthompson5910 (jim_thompson5910):

In this case, B = pi/3

jimthompson5910 (jim_thompson5910):

T = ??

OpenStudy (anonymous):

2pi/(pi/3)?

OpenStudy (anonymous):

so..six?

jimthompson5910 (jim_thompson5910):

T = 6, yep

jimthompson5910 (jim_thompson5910):

The frequency f is found by this formula f = 1/T

OpenStudy (anonymous):

so the answer is 1/6?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

thanks a ton! a lot of stuff just cleared up!

jimthompson5910 (jim_thompson5910):

I'm glad it has

OpenStudy (anonymous):

sooo, for d = 9 cos ((pi/2)t), would the period be 4? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

thanks :)

jimthompson5910 (jim_thompson5910):

np